OpenStudy (kj4uts):

What is the value of the fourth term in a geometric sequence for which a1=10 and r=0.5?

OpenStudy (tkhunny):

a1 = 10 a2 = 10*0.5 a3 = 10*0.5*0.5 etc.

OpenStudy (imstuck):

The formula for that is this:\[a _{n}=a _{1}* r ^{n-1}\]Your first term is a1 and the n is 4 because that is the number you are looking for...the fourth term.